# Information Seeking Using AI Assistants

### v1

@@ VI Threats to Validity
External. Our sample was recruited through LinkedIn, X, and internal mailing lists, which might not be representative of the broader developer community. Developers who are not active on these platforms or in these networks might have different experiences with AI tools. To mitigate this, we also engaged with local developer communities and recruited participants through snowball sampling. Human-centered empirical studies are also prone to generalizability issues, due in part to concerns like sample size. However, the goal of our study is not to be generalizable, but rather transferable [44]. We ensure this by using a mixed-method approach where we supplement our survey findings with interviews. In our qualitative analysis, we followed best practices to ensure rigor. We did not report qualitative data using quantitative methods to prevent misinterpretation of our findings [45]. To further ensure the validity of our thematic analysis, we invited an external auditor to review our methodology and findings, providing an additional layer of validation.

### v2

@@ 6. Threats to Validity
External. Our sample was recruited through LinkedIn, X, and internal mailing lists, which might not be representative of the broader developer community. Developers who are not active on these platforms or in these networks might have different experiences with AI tools. To mitigate this, we also engaged with local developer communities and recruited participants through snowball sampling. Human-centered empirical studies are also prone to generalizability issues, due in part to concerns like sample size. However, the goal of our study is not to be generalizable, but rather transferable (Daniel, 2019). We ensure this by using a mixed-method approach where we supplement our survey findings with interviews. In our qualitative analysis, we followed best practices to ensure rigor. We did not report qualitative data using quantitative methods to prevent misinterpretation of our findings (Maxwell, 2010). To further ensure the validity of our thematic analysis, we invited an external auditor to review our methodology and findings.